# How to Train Your LLM Web Agent: A Statistical Diagnosis

### v1

@@ Abstract
LLM-based web agents have recently made significant progress, but much of it has occurred in closed-source systems—widening the gap with open-source alternatives. Progress has been held back by two key challenges—first, a narrow focus on single-step tasks that overlooks the complexity of multi-step web interactions, and second, the high compute costs required to post-train LLM-based web agents. To address this, we present the first statistically grounded study on compute allocation for LLM web-agent post-training. Our approach uses a two-stage pipeline, training a Llama 3.1 8B student to imitate a Llama 3.3 70B teacher via SFT, followed by on-policy reinforcement learning. We find this process highly sensitive to hyperparameter choices — exhaustive sweeps are impractical. To spare others from expensive trial-and-error, we sample 1,370 configurations and use bootstrapping to estimate effective hyperparameters. Our results show that combining SFT with on-policy RL consistently outperforms either approach alone on both WorkArena and MiniWob++. Further, this strategy only requires 55% of the compute to match the peak of pure SFT on MiniWob++, pushing the compute–performance Pareto frontier and is the only strategy that can close the gap with closed-source models.
@@
Figure 1: Compute–performance frontier on MiniWoB++ (results averaged over two seeds). The blue curve shows pure SFT on teacher demonstrations. Warm-colored curves represent hybrid runs that branch off from SFT checkpoints and continue training with RL. Early transitions to RL push the Pareto frontier—achieving higher success rates for the same compute—and is the only approach able to achieve over $30\%$ improvement on both held-out goals (left) and held-out tasks (right) closing the gap between open and closed-source models.
@@ 1 Introduction
Large language model (LLM) agents for web interfaces have advanced rapidly, but open-source systems still trail proprietary ones. Bridging this gap would allow organizations to train smaller, cost-efficient agents tailored to their needs while maintaining data privacy. Yet, despite impressive progress of open-source models in domains like math and code generation, advances in training web-capable LLM agents remain limited by two persistent challenges: a lack of attention to multi-turn, long-horizon tasks, and the high cost and low reproducibility of current training pipelines.
Most research centers on single-step tasks like code or math—domains with rapid feedback and simplified credit assignment—which fall short of real-world web environments requiring sequential decisions and long-horizon planning. Recent benchmarks like WebArena [25], WorkArena [10], OSWorld [22], and The Agent Company [23] have exposed how brittle current methods become under delayed rewards, sparse feedback, and compounding errors. Addressing these settings demands not just better agents, but reproducible, compute-efficient training pipelines—an area we directly tackle.
However, building such pipelines is nontrivial. Modern LLM post-training often involves a combination of Supervised Fine-Tuning (SFT) and Reinforcement Learning (RL), with performance sensitive to a large number of interacting hyperparameters [11]. Single-seed results are noisy and misleading, but exhaustive tuning and multi-seed sweeps, e.g. [2], remain infeasible for most labs due to the cost of LLM training. This makes it all the more critical to design pipelines that are not only effective but statistically robust and accessible within realistic compute budgets.
In this paper, we tackle these gaps by providing a statistically driven diagnosis of training LLM agents for web-based, multi-step tasks. Specifically, we study how to allocate compute between expensive, high-quality off-policy demonstrations from a large teacher model and cheaper on-policy rollouts from a smaller student model. We analyze this tradeoff across two levels of generalization: held-out goals—tasks encountered during training but with novel goals—and held-out tasks, which are entirely unseen during training.
@@
Putting things together, we show how our study yields several actionable insights. First, branching into RL early—but not immediately—after SFT leads to better outcomes. This hybrid strategy consistently outperforms pure SFT and pure RL, and reaches the maximal performance of pure SFT while requiring only 55% of the compute, effectively pushing the compute–performance Pareto frontier. It is also the only strategy that can close the gap with closed-source models. Second, curriculum learning and error log feedback help the less SFT warmup has been applied but can become counterproductive thereafter. Third, in GRPO, applying zero-advantage filtering consistently improve performance, while dividing the advantage with the standard deviation and wether to use an importance ratio are dependent on the amount of SFT warmup done. Fourth, decoding temperature is consistently critical, while learning rate and discount rate must also be carefully tuned.
@@ 2 Background
This section consolidates the algorithmic ingredients used throughout the paper: (i) the MDP formulation of web-based language agents, (ii) SFT on expert traces, (iii) GRPO for RL, and (iv) curriculum and normalization techniques that stabilize training.
@@ 2.1 Language Agents as MDPs
We model each task as a Markov Decision Process (MDP) $\mathcal{M}=\langle\mathcal{S},\mathcal{A},P,r,\rho_{0},\gamma\rangle$. A state $\mathbf{s}_{t}\in\mathcal{S}$ is a textual context–in our case a prompt and an action $a_{t}\in\mathcal{A}$ is a textual response generated by the agent. Each action $a_{t}$ consists of a sequence of tokens $o_{1_{t}:K_{t}}$ sampled autoregressively from the policy $\pi_{\theta}(a_{t}|\mathbf{s}_{t})$ parameterized by an LLM sampled with temperature $\rho_{\text{LLM}}$. The environment then returns a scalar reward $r_{t}\in\{-1,1\}$ indicating task failure/success. In our setting we assume the environment dynamics are $p(\mathbf{s}_{t+1}\mid a_{t})\in P$ We optimize the policy $\pi_{\theta}$ to maximize the expected discounted return:
@@
2.2 Off-policy boot-strapping via SFT
We first imitate a stronger expert policy $\pi_{E}$ by minimizing the cross-entropy loss
@@ 2.2 Off-policy boot-strapping via SFT
SFT offers a high-quality, low-variance gradient but is inherently off-policy which can lead to poor generalization [6].
@@ 2.3 On-Policy Improvement with Multi-Turn GRPO
After the initial SFT warmup phase, we continue training using on-policy RL using GRPO as the optimization algorithm. Like REINFORCE [18], GRPO maximizes the expected return using a policy gradient objective. However, GRPO introduces additional structure by leveraging per-goal advantage normalization and importance weighting, wherein each trajectory is associated with a specific goal—in our context, referring to the seed of the target task. For a given goal $g$, the group-normalized advantage function is:
@@
where $\mathbf{R}_{t,g}=(r_{t,1},\dots,r_{t,G})$ is the set of rewards across the $G$ goals at timestamp $t$. In addition to this–similar to proximal policy optimization (PPO) [2]–an importance-ratio is often applied to the GRPO objective:
@@
Zero-advantage filtering.
@@ 3 Methodology
Our training pipeline consists of two sequential stages—SFT followed by RL—framed as a resource allocation problem. We also describe our hyperparameter sweep strategy and statistical analysis protocol that consolidates hundreds of runs into reliable conclusions. We evaluate our recipe along two axes: compute cost, measured in FLOPs (using the formula from [1]), and model performance, assessed on both unseen training goals and held-out testing tasks.
@@ 3.1 Estimating the Uncertainty of the Hyperparameter Selection Process
Across the different SFT checkpoints, we sample 1,370 distinct configurations with ten varying hyperparameters (see Appendix F for details). Our objective is to study the effect of various hyperparameters (HP) on the downstream success rate of our trained agents. This comes with two important considerations. First, if we change the value of, e.g., the batch size, and we want to know if a bigger batch size is better, the learning rate and other parameters need to be readjusted close to their optimal configuration (under a fixed budget). Secondly, to account for noise, we would need to restart the same experiment several times to avoid spurious conclusions. In practice, this is out of reach. For a more computationally friendly approach, we resort to bootstrapping the collection of trials over different hyperparameter configurations.
@@
From the full set of 1,370 training runs, we perform bootstrap resampling by drawing individual runs (with replacement). For each resample, we identify the best-performing configuration and repeat this process 1,000 times. We also compute the fraction of times each hyperparameter value "wins", which serves as an estimated probability that it belongs to the global optimum. This procedure serves two purposes: to estimate the maximum relative improvement a specific hyper-parameter provides—while accounting for variation in other parameters—and to offer uncertainty estimates in the form of win-rate distributions between different configurations. In addition, to better understand how optimal hyper-parameters may change depending on the amount of SFT warmup we apply this analysis across various SFT checkpoints.
@@
We provide a detailed explanation of this procedure in Algorithm 1. Additionally in Appendix F we describe how use this procedure to obtain the results reported in Figure 1.
@@
2:$\mathcal{W}=\texttt{dict}(\ )$triangleright Win counts per HP value
@@
4:$H_{\text{vals}}=\texttt{unique}(\texttt{Results}[H])$triangleright Distinct values of the hyperparameter
@@
8:     $h^{*}=\operatorname{argmax}_{h}\texttt{sample}_{b}[M]$triangleright Get the winning HP value for $\texttt{sample}_{b}$
@@ 4 Experimental Setup
We evaluate our approach using Llama 3.3 70B as the expert model to generate demonstration traces, and Llama 3.1 8B as the student model for fine-tuning. Both models operate with a 16k token context window to handle the complexity of web-based tasks.
@@
Our experiments focus on two benchmarks. The first is MiniWoB++, a suite of 30 medium-horizon web interaction tasks, where we observe that optimal policies typically complete tasks in 2 to 5 steps. The second is WorkArena [10], a more challenging benchmark of 33 enterprise knowledge-work tasks, where we observe optimal policies generally complete tasks in 3 to 10 steps. These benchmarks provide a representative spectrum of sequential decision-making challenges faced by interactive LLM agents. Both benchmarks are depicted in Figure 3.
@@
MiniWoB++ provides raw HTML trees, whereas WorkArena leverages accessibility trees (AxTrees), which we truncate to 16k tokens to meet hardware constraints. The agent operates in a discrete action space composed of high-level UI primitives: noop, fill(node, text), click(node), type(node, text), select_option(node, option), scroll(node), and hover(node). This abstraction allows the agent to interact effectively with diverse web interfaces. All agents employ chain-of-thought prompting [21]. We also experiment with error log feedback, allowing the agent to receive explicit error messages when it takes invalid actions.
@@
The environments in this study are stochastic; each task is initialized with a random seed. We define a “goal” as a specific random seed for a given task. The evaluation protocol assesses generalization at two levels: (i) performance on held-out goals within the training tasks, and (ii) performance on entirely new, unseen held-out tasks. We consider both settings important as they capture different aspects of generalization—held-out goals test reliability on trained tasks, while held-out tasks test the model’s ability to transfer skills learnt in training to entirely new situations.
@@
We track the total floating-point operations (FLOPs) consumed during both SFT and GRPO phases, following the procedure in Section 3 and Appendix B. For RL branching, we start from the SFT run that (i) attains the highest SFT performance and (ii) exhibits stable learning, allowing us to sample checkpoints at regular intervals and cleanly study the effect of SFT warm-up. Because this run sits in the top 10% of all SFT trials by area-under-the-training-curve (AUC)—the mean success rate across all checkpoints of the run, serving as a proxy for overall training efficiency, we match that selection pressure for RL by averaging the top two seeds out of ten RL trials, thus approximating the 90th-percentile of the RL distribution and ensuring a fair compute-aware comparison between strategies.
@@ 5 Main Results and Compute Trade-Offs
LABEL:tab:performance_comparison summarizes the results on MiniWoB++ and WorkArena. Combining SFT with RL consistently yields the best performance among student models. Pure SFT and pure RL each fall short, with RL from scratch particularly struggling due to sparse rewards and unstable learning dynamics. These results highlight the complementary strengths of expert demonstrations and on-policy fine-tuning.
@@
Our analysis focuses on the trade-off between costly but high-quality teacher demonstrations and cheaper, noisier on-policy rollouts. To examine this, we branch RL fine-tuning from SFT checkpoints sampled at fixed intervals along the training trajectory. Figure 1 illustrates the compute–performance frontier on MiniWoB++, where runs that combine SFT with RL (warm colors) consistently outperform pure SFT (blue), pushing the Pareto frontier forward. Early branching into RL unlocks substantial gains in compute efficiency, delivering stronger performance at lower compute budgets. For example, SFT+RL reaches the maximal performance of pure SFT on the test set (achieved at approximately 11 exaFLOPs with pure SFT) using only around 6 exaFLOPs—representing a compute saving of roughly 45% (11 vs 6 exaFLOPs) a compute saving of roughly 45%.
Notably, this is the only strategy that closes the performance gap with closed-source models like GPT-4o. The trend is consistent across both held-out goals and held-out tasks: warm-started RL yields higher performance than either SFT or RL alone, reinforcing the importance of blending expert supervision with on-policy learning. These findings underscore the need to balance sample efficiency from demonstrations with the compute efficiency of on-policy learning—a key consideration for scaling LLM-agent training.
@@
Despite extensive post-training, agent performance on the WorkArena benchmark plateaus at around 40% after just 9–10 epochs. This stagnation appears to stem from the intrinsic difficulty of certain tasks—such as sorting and filtering—which even the Llama 3.3 70B teacher model struggles to solve (see Figure 3). A per-task breakdown shows that while both SFT and RL agents gradually close the performance gap with the teacher model, with RL achieving a slightly higher final success rate, a significant portion of tasks (14 out of 33) remain completely unsolved. These failures are attributed to either the limitations of the teacher model or the sparsity of reward signals, both of which hamper the learning process. On-policy RL exploration proves ineffective in overcoming these challenges due to the lack of foundational skills and informative feedback. These findings underscore the need for more effective methods to address complex tasks under sparse reward settings. Additional per-task performance results for WorkArena and Miniwob are provided in Appendix A.
@@
During extended training we found that giving agents full administrative privileges invites creative—but problematic—short-cuts. A striking example comes from the WorkArena “Order hardware devices” task: rather than navigate through the catalogue as intended, the agent simply edited its homepage to add a direct link to the ordering form. While this hack fulfils the success condition, it violates the task’s spirit and, worse, the modifications persist across sessions. Other agents entering the environment later inherit the altered UI, leading to failures unrelated to their own policies.
In general, agents exploited admin rights to create or delete elements, reshaping pages in ways that saved clicks but destabilised the environment. These findings argue for a tighter sandbox—enough access for exploration and accomplishment, but safeguards against permanent, instance-wide side-effects.

### v3

@@ Abstract
LLM-based web agents have recently made significant progress, but much of it has occurred in closed-source systems, widening the gap with open-source alternatives. Progress has been held back by two key challenges, first, a narrow focus on single-step tasks that overlooks the complexity of multi-step web interactions, and second, the high compute costs required to post-train LLM-based web agents. To address this, we present the first statistically grounded study on compute allocation for LLM web-agent post-training. Our approach uses a two-stage pipeline, training a Llama 3.1 8B student to imitate a Llama 3.3 70B teacher via SFT, followed by on-policy reinforcement learning. We find this process highly sensitive to hyperparameter choices in setting where exhaustive sweeps are impractical. To spare others from expensive trial-and-error, we sample 1,370 configurations and use bootstrapping to estimate effective hyperparameters. Our results show that combining SFT with on-policy RL consistently outperforms either approach alone on both WorkArena and MiniWob++. Further, this strategy only requires 55% of the compute to match the peak of pure SFT on MiniWob++, pushing the compute–performance Pareto frontier and is the only strategy that can close the gap with closed-source models.
@@
Figure 1: Compute–performance frontier on MiniWoB++ (results averaged over two seeds). The blue curve shows pure SFT on teacher demonstrations. Warm-colored curves represent hybrid runs that branch off from SFT checkpoints and continue training with RL. Early transitions to RL push the Pareto frontier achieving higher success rates for the same compute and is the only approach able to achieve over $30\%$ improvement on both held-out goals (left) and held-out tasks (right) closing the gap between open and closed-source models. See fig.˜4 for the corresponding plot with Qwen2.5 7B.
@@ 1 Introduction
Large language model (LLM) agents for web interfaces have advanced rapidly, but open-source systems still trail proprietary ones. Bridging this gap would allow organizations to train smaller, cost-efficient agents tailored to their needs while maintaining data privacy. Yet, despite impressive progress of open-source models in domains like math and code generation, advances in training web-capable LLM agents remain limited by mainly by the lack of attention to multi-turn, long-horizon tasks and the high cost and low reproducibility of current training pipelines.
Most research centers on single-step tasks like code or math domains with rapid feedback and simplified credit assignment which fall short of real-world web environments requiring sequential decisions and long-horizon planning. Recent benchmarks like WebArena [29], WorkArena [10], OSWorld [26], and The Agent Company [27] have exposed how brittle current methods become under delayed rewards, sparse feedback, and compounding errors. Addressing these settings demands not just better agents, but reproducible, compute-efficient training pipelines an area we directly tackle.
However, building such pipelines is nontrivial. Modern LLM post-training often involves a combination of Supervised Fine-Tuning (SFT) and Reinforcement Learning (RL), with performance sensitive to a large number of interacting hyperparameters [13]. Single-seed results are noisy and misleading, but exhaustive tuning and multi-seed sweeps (e.g. [2]) remain infeasible for most labs due to the cost of LLM training. This makes it all the more critical to design pipelines that are not only effective but statistically robust and accessible within realistic compute budgets.
In this paper, we tackle these gaps by providing a statistically driven diagnosis of training LLM agents for web-based, multi-step tasks. Specifically, we study how to allocate compute between expensive, high-quality off-policy demonstrations from a large teacher model and cheaper on-policy rollouts from a smaller student model. We analyze this tradeoff across two levels of generalization, held-out goals being tasks encountered during training but with novel goals and held-out tasks, which are entirely unseen during training.
@@
Putting things together, we show how our study yields several actionable insights. First, branching into RL early, but not immediately after SFT leads to better outcomes. This hybrid strategy consistently outperforms pure SFT and pure RL, and reaches the maximal performance of pure SFT while requiring only 55% of the compute, effectively pushing the compute–performance Pareto frontier. It is also the only strategy that can close the gap with closed-source models. Second, curriculum learning and error log feedback help the less SFT warmup has been applied but can become counterproductive thereafter. Third, in GRPO, applying zero-advantage filtering consistently improve performance, while dividing the advantage with the standard deviation and wether to use an importance ratio are dependent on the amount of SFT warmup done. Fourth, decoding temperature is consistently critical, while learning rate and discount rate must also be carefully tuned.
@@ 2 Background
This section consolidates the algorithmic ingredients used throughout the paper: (i) the MDP formulation of web‑based language agents, (ii) SFT on expert traces, (iii) GRPO for RL, and (iv) curriculum and normalization techniques that stabilize training.
@@ 2.1 Language Agents as MDPs
We model each task as a Markov Decision Process (MDP) $\mathcal{M}=\langle\mathcal{S},\mathcal{A},P,r,\rho_{0},\gamma\rangle$. A state $\mathbf{s}_{t}\in\mathcal{S}$ is a textual context–in our case a prompt and an action $a_{t}\in\mathcal{A}$ is a textual response generated by the agent. Each action $a_{t}$ consists of a sequence of tokens $o_{1_{t}:K_{t}}$ sampled autoregressively from the policy $\pi_{\theta}(a_{t}|\mathbf{s}_{t})$ parameterized by an LLM sampled with temperature $\rho_{\text{LLM}}$. The environment then returns a scalar reward $r_{t}\in\{-1,1\}$ indicating task failure/success. In our setting we assume the environment dynamics are $p(\mathbf{s}_{t+1}\mid a_{t})\in P$ . We optimize the policy $\pi_{\theta}$ to maximize the expected discounted return:
@@
2.2 Off‑policy boot‑strapping via SFT
We first imitate a stronger expert policy $\pi_{E}$ by minimizing the cross‑entropy loss
@@ 2.2 Off‑policy boot‑strapping via SFT
SFT offers a high‑quality, low‑variance gradient but is inherently off‑policy which can lead to poor generalization [6].
2.3 Rejection Fine-Tuning
Similar to SFT, Rejection Fine-Tuning (RFT) optimizes the policy using successful trajectories, but in this case bootstrapped of the training policy rather than a larger expert
$\mathcal{L}_{\mathrm{RFT}}(\theta)=-\mathbb{E}_{\tau\sim\pi_{\theta}}\Big[\mathbf{1}_{\{r_{t}=1\}}\sum_{t=0}^{T}\log\pi_{\theta}(a_{t}\mid s_{t})\Big].$ (3)
@@ 2.4 On-Policy Improvement with Multi-Turn GRPO
After the initial SFT warmup phase, we continue training using on-policy RL using GRPO as the optimization algorithm. Like REINFORCE [21], GRPO maximizes the expected return using a policy gradient objective. However, GRPO introduces additional structure by leveraging per-goal advantage normalization and importance weighting, wherein each trajectory is associated with a specific goal in our context, referring to the seed of the target task. For a given goal $g$, the group-normalized advantage function is:
@@
where $\mathbf{R}_{t,g}=(r_{t,1},\dots,r_{t,G})$ is the set of rewards across the $G$ goals at timestamp $t$. In addition to this similar to proximal policy optimization (PPO) [2] an importance-ratio is often applied to the GRPO objective:
@@
Zero‑advantage filtering.
@@ 3 Methodology
Our training pipeline consists of two sequential stages SFT followed by RL framed as a resource allocation problem. We also describe our hyperparameter sweep strategy and statistical analysis protocol that consolidates hundreds of runs into reliable conclusions. We evaluate our recipe along two axes: compute cost, measured in FLOPs (using the formula from [1]), and model performance, assessed on both unseen training goals and held-out testing tasks.
@@ 3.1 Estimating the Uncertainty of the Hyperparameter Selection Process
Across the different SFT checkpoints, we sample 1,370 distinct configurations with ten varying hyperparameters (see Appendix˜L for details). Our objective is to study the effect of various hyperparameters (HP) on the downstream success rate of our trained agents. This comes with two important considerations. First, if we change the value of, e.g., the batch size, and we want to know if a bigger batch size is better, the learning rate and other parameters need to be readjusted close to their optimal configuration (under a fixed budget). Secondly, to account for noise, we would need to restart the same experiment several times to avoid spurious conclusions. In practice, this is out of reach. For a more computationally friendly approach, we resort to bootstrapping the collection of trials over different HP configurations. The boostrap algorithm has many desriable properties such as being a consistent and unbiased estimator for the sampling distribution of any data statistic without requiring parametric assumptions (see Appendix˜K for details), thus allowing us to reliably obtain robust estimates of both the expected value and variance of the success rate given by any HP.
@@
From the full set of 1,370 training runs, we perform bootstrap resampling by drawing individual runs (with replacement). For each resample, we identify the best-performing configuration and repeat this process 1,000 times. We also compute the fraction of times each hyperparameter value "wins", which serves as an estimated probability that it belongs to the global optimum. This procedure serves two purposes: to estimate the maximum relative improvement a specific hyper-parameter provides while accounting for variation in other parameters and to offer uncertainty estimates in the form of win-rate distributions between different configurations. In addition, to better understand how optimal hyper-parameters may change depending on the amount of SFT warmup we apply this analysis across various SFT checkpoints.
@@
We provide a detailed explanation of this procedure in Algorithm 1. Additionally in Appendix˜J we describe how use this procedure to obtain the results reported in Figure˜1.
@@
[…]
@@
8:  $h^{*}=\operatorname{argmax}_{h}\texttt{sample}_{b}[M]$ $\triangleright$ Get the winning HP value for $\texttt{sample}_{b}$
@@ 4 Experimental Setup
We evaluate our approach using two teacher models, Llama 3.3 70B and Qwen 2.5 72B to generate SFT traces, with Llama 3.1 8B and Qwen 2.5 7B acting as the student model for fine-tuning. All models operate with a 16k token context window to handle the complexity of web-based tasks. For the LLama models we report results using GRPO, while for Qwen we report results using RFT as we found it more stable during training.
@@
Our experiments focus on two benchmarks. The first is MiniWoB++, a suite of 30 medium-horizon web interaction tasks, where we observe that optimal policies typically complete tasks in 2 to 5 steps. The second is WorkArena [10], a more challenging benchmark of 33 enterprise knowledge-work tasks, where we observe optimal policies generally complete tasks in 3 to 10 steps. These benchmarks provide a representative spectrum of sequential decision-making challenges faced by interactive LLM agents. Both benchmarks are depicted in Figure˜7.
@@
MiniWoB++ provides raw HTML trees, whereas WorkArena leverages accessibility trees (AxTrees), which we truncate to 16k tokens to meet hardware constraints. The agent operates in a discrete action space composed of high-level UI primitives: noop, fill(node, text), click(node), type(node, text), select_option(node, option), scroll(node), and hover(node). This abstraction allows the agent to interact effectively with diverse web interfaces. All agents employ chain-of-thought prompting [25]. We also experiment with applying error log feedback, allowing the agent to receive explicit error messages when it takes invalid actions (during both training and testing).
@@
Both environments are stochastic, and we control for randomness during training by fixing the task seed. Each task–seed pair is referred to as a ”goal”. The evaluation protocol assesses generalization at two levels: (i) performance on held-out goals within the training tasks, and (ii) performance on entirely new, unseen held-out tasks. We consider both settings important as they capture different aspects of generalization held-out goals test reliability on trained tasks, while held-out tasks test the model’s ability to transfer skills learnt in training to entirely new situations.
@@
We track the total floating‑point operations (FLOPs) consumed during both SFT and GRPO phases, following the procedure in Section 3 and Appendix D. For RL branching, we start from the SFT run that (i) attains the highest SFT performance and (ii) exhibits stable learning, allowing us to sample checkpoints at regular intervals and cleanly study the effect of SFT warm‑up. Because this run sits in the top 10% of all SFT trials by area‑under‑the‑training‑curve (AUC) the mean success rate across all checkpoints of the run, serving as a proxy for overall training efficiency, we match that selection pressure for RL by averaging the top two seeds out of ten RL trials, thus approximating the 90th‑percentile of the RL distribution and ensuring a fair compute‑aware comparison between strategies.
@@ 5 Main Results and Compute Trade-Offs
Table˜1 summarizes the results on MiniWoB++ and WorkArena. Combining SFT with RL consistently yields the best performance among student models. Pure SFT and pure RL each fall short, with RL from scratch particularly struggling due to sparse rewards and unstable learning dynamics. These results highlight the complementary strengths of expert demonstrations and on-policy fine-tuning.
@@
Table 1: Comparison of our method with baselines on the held-out train and test splits of MiniWoB++ and WorkArena. Here $\pm$ are the averaged standard errors computed separately for each of the two top seeds.
Model MiniWoB++ WorkArena
Held-out Goals Held-out Tasks Held-out Goals Held-out Tasks
[…]
@@
Our analysis focuses on the trade-off between costly but high-quality teacher demonstrations and cheaper, noisier on-policy rollouts. To examine this, we branch RL fine-tuning from SFT checkpoints sampled at fixed intervals along the training trajectory. Figure 1 illustrates the compute–performance frontier on MiniWoB++, where runs that combine SFT with RL (warm colors) consistently outperform pure SFT (blue), pushing the Pareto frontier forward. Early branching into RL unlocks substantial gains in compute efficiency, delivering stronger performance at lower compute budgets. For example, SFT+RL reaches the maximal performance of pure SFT on the test set (achieved at approximately 11 exaFLOPs with pure SFT) using only around 6 exaFLOPs translating to roughly 45% of compute saved (11 vs 6 exaFLOPs).
Notably, this is the only strategy that closes the performance gap with closed-source models like GPT-4o. The trend is consistent across both held-out goals and held-out tasks: warm-started RL yields higher performance than either SFT or RL alone, reinforcing the importance of blending expert supervision with on-policy learning. These findings underscore the need to balance sample efficiency from demonstrations with the compute efficiency of on-policy learning, and the results generalize to Qwen2.5 7B (see fig.˜4).
@@
Despite extensive post-training, agent performance on the WorkArena benchmark plateaus at around 40% after just 9–10 epochs. This stagnation appears to stem from the intrinsic difficulty of certain tasks (e.g as sorting and filtering tasks) that even frontier models struggles to solve (see Figure˜2). A per-task breakdown shows that while both SFT and RL agents gradually close the performance gap with the teacher model, with RL achieving a slightly higher final success rate, a significant portion of tasks (14 out of 33) remain completely unsolved. These failures are attributed to either the limitations of the teacher model or the sparsity of reward signals, both of which hamper the learning process. On-policy RL exploration proves ineffective in overcoming these challenges due to the lack of foundational skills and informative feedback. These findings underscore the need for more effective methods to address complex tasks under sparse reward settings. Additional per-task performance results for WorkArena and Miniwob are provided in Appendix˜C.
@@
During extended training we found that giving agents full administrative privileges invites creative but problematic shortcuts. For instance, in WorkArena’s “Order hardware devices” task, rather than navigating through the catalogue as intended, the agent simply edits its homepage to add a direct link to the ordering form. While this hack fulfills the success condition, it violates the task’s spirit and, worse, the modifications persist across sessions. Other agents entering the environment later inherit the altered UI, leading to failures unrelated to their own policies.
In general, agents exploited admin rights to create or delete elements, reshaping pages in ways that saved clicks but destabilised the environment. These findings argue for a tighter sandbox enough access for exploration and accomplishment, but safeguards against permanent, instance‑wide side‑effects.